The nth term of a G.P. is 2x3^n for all n. Find (a) the first term (b) the common ratio of the G.P.
Answer by mananth(16946)


2 * 3^n
for n=1 t1= 6
for n=2 t2= 18
for n=3 . t3 = 54
common ratio = 18/6 = 54/18 =3